A single layer of No. 24 AWG (dia=0.0201 in) commercial iron wire is wound over a ceramic tube whose diameter is 3.25 in. if the total wire resistance is 41 ohms, determine the number of turns.
1
Expert's answer
2020-09-15T10:04:22-0400

Solution

Given data

Diameter (d)=0.0201 in

Resistance (R)=41"\Omega"

Circumference of ceramic tube(c)"=\pi D"

="\pi\frac{3.25}{12}=0.85ft" =1turn


And CM="\frac{0.021}{0.001}=404.01mil^2"

Resistance

"R=\rho \frac{L}{CM}"


"L=R\frac{CM}{\rho}=\frac{41\times404.01}{75}=221ft"


No. of turns

"N=\frac{L}{C}=\frac{221}{0.85}=259.83"


"N=260"

Therefore number of turns are 260.